Write all the factors of the following numbers: 
24 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

24 = 1 × 24 24 = 2 x 12 24 = 3 x 8
24 = 4 × 6 24 = 6 x 4
Stop here since 4 and 6 have occurred earlier
Hence, the factors of 24 are 1, 2, 3, 4, 6, 8, 12 and 24

​Thanks